# Forge: Foundational Optimization Representations from Graph Embeddings

### v1

@@ 1 Introduction
With this vision in mind, we propose Forge; Foundational Optimization Representations from Graph Embeddings. Forge is a foundational model to generate useful MIP embeddings through a pre-training framework that learns structural representations at the instance level from a broad distribution of MIP instances without requiring access to their solutions.
@@
  • For the unsupervised setting, we demonstrate that our instance level MIP embeddings clusters previously unseen instances across multiple problem distributions with high accuracy.
  • For the supervised setting, we demonstrate that pre-trained Forge embeddings can be further fine-tuned for various downstream tasks using a small number of labeled data. We evaluate Forge on two radically different tasks to showcase its versatility: predicting variables for warm-starts and predicting integrality gap for cut-generation. Notably, a single pre-trained Forge model is fine-tuned and applied across diverse problem types including Combinatorial Auctions (CA), Set Covering (SC), Generalized Independent Set (GIS), and Minimum Vertex Cover (MVC).
@@
The remainder of the paper is structured as follows: we start with a brief background of CO in Section §2 followed by the methodology for generating MIP embeddings in Section §3. Sections §4 and §5 show experiments to demonstrate the efficacy of Forge in unsupervised and supervised settings respectively.
@@ 3 Forge: Unsupervised Approach to Learn MIP Embeddings
It is common to represent MIP instances as a bipartite graph between variables and constraints augmented with node features as pioneered by (Gasse et al., 2019). This is typically followed by training a GNN in a supervised fashion for a specific downstream task on a certain problem class. For example, in Han et al. (2023), the GNN is used for predicting variables for warm-starting a MIP solver trained on Set Cover (SC) and Independent Set (IS) problems. Similarly, in Li et al. (2025), a GNN is used for predicting the integrality gap. There are many other examples of this architecture in the literature as covered in our related works section (§6). Notice that, all of these methods require supervision and do not yield a general-purpose MIP embeddings at the instance level. Taking this a step further; our goal is to learn the structure of MIP instances in an unsupervised manner.
@@
A) MIP-to-BP: Given a MIP instance, we start with its bipartite (BP) representation and node features (Gasse et al., 2019; Ferber et al., 2022; Yau et al., 2024; Chen et al., 2023).
Each node in this bipartite graph represents either a constraint or a variable, with edges indicating which variables are part of which constraints. Each node is also associated with node features and each edge is weighted by the coefficient of the variable in the constraint. Traditionally, node features depend on extracting information from the internal branch-and-bound tree search when solving an instance. For example, there are 18 input node features used in Gasse et al. (2019). We do not attempt to solve or depend on the solution of the instance. Instead, Forge only uses the basic properties of the input instance. For each constraint node, we introduce 4 features composed of its sense (i.e., >>, << or ==) and the RHS value. For each variable node, we introduce 6 features composed of its type (integer, binary, continuous), upper/lower bound, and the coefficient in the objective function. In total, for each node, we obtain a vector of size 10, padded with zeros accordingly based on node type (Figure 1(A)).
@@
C) Vector Quantized codebook: To preserve global structure, we introduce a vector quantized codebook with $k$ discrete codes. These discrete codes act as a “vocabulary”, akin to language models, across MIP instances of various domains and difficulties, thereby preserving global structure. The design follows the approaches developed in computer vision (Van Den Oord et al., 2017; Yu et al., 2022; Lee et al., 2022) and their structure-aware graph tokenizer extension as proposed by Yang et al. (2024) (Figure 1(C)).
D) GNN-to-CW: GNN embeddings are passed into a vector quantizer which consists of a codebook with $k$ discrete codes. The codebook maps each variable and constraint node to a discrete code. Each code is then mapped into a $d$ dimensional codeword (CW), yielding constraint and variable CW representations. These representations are the same dimension as the hidden GNN layers (Figure 1(D)).
@@
E) Loss Function: The overall loss function minimizes the edge reconstruction loss, the node feature reconstruction loss as well as losses related to the vector quantization. Specifically, the loss function is given by:
@@
Intuitively, the codebook loss in Eq. 3 can be interpreted as k-means clustering, where the codewords $cw_{i}$ (akin to cluster centroids) are updated to move closer to the node embeddings $h_{i}$ and the node embeddings $h_{i}$ are fixed in place due to the stop-gradient operator. Conversely, the commitment loss in Eq. 4 fixes the codewords $cw_{i}$ using the stop-gradient operator, and instead, updates the embeddings $h_{i}$ to move towards the codewords. The hyperparameter $\alpha$ weighs the importance of the commitment loss.
@@
Local representations, where each node in the bipartite graph of the MIP instance is assigned a discrete code. Each code is further mapped to a codeword which becomes the variable and/or constraint embedding.
Global representations, where we leverage the distribution of codes for the given MIP instance. As shown in Figure 2, each MIP instance can be represented with an embedding of size $vocabulary=\lvert codebook\rvert$ whereby each value denotes the frequency of codes present in the instance. Figure 2 shows a codebook of size 5 (for brevity) and the resulting MIP embedding $\vec{emb}=[3,2,3,2,0]$ from the bipartite graph.
@@ 4 Initial Analysis of Forge Embeddings
Figure 3: 2-D PaCMAP visualization of clustering D-MIPLIB instances using MIP instance embeddings from (a) the mean readout of node embeddings from the GNN layer in Forge, (b) the two-hop label propagation of input node features, and (c) the distribution of discrete codes produced by Forge. Normalized Mutual Information (NMI), averaged over 10 runs of $k$-means clustering, is reported for each method.
@@
Training: We use 600 instances from MIPLIB, sorted by size to ensure the resulting bipartite graphs fit on GPU memory. We further generate two more instances from each MIPLIB instance by randomly dropping 5% and 10% of constraints. Note that dropping constraints does not impact feasibility of the instance. In total, we obtain 1,800 MIP instances to train Forge in an unsupervised fashion with reconstruction loss as described earlier in (§3). We use two GraphSage layers with 1024 dimensions for the hidden GNN layers ($d$) and a codebookcodebook with 5000 codes (i.e., the size of the vocabulary).
@@
Testing: We evaluate the trained Forge model on 1,050 instances from D-MIPLIB organized into 21 problem type-difficulty pairs. These are distributed as Set Cover (easy, medium, hard), Maximum Independent Set (easy, medium), Minimum Vertex Cover (easy, medium, hard), Generalized Independent Set (easy, medium, hard, very-hard, very-hard2, ext-hard), Combinatorial Auction (very-easy, easy, medium, very-hard, very-hard2), Item Placement (very-hard) and Maritime Inventory Routing Problem (medium), thereby covering a broad spectrum of problem types, sizes and difficulty levels.
@@
MIP Embeddings: These 1,050 instances are passed through Forge to generate one embedding vector per MIP instance. This featurization uses the distribution of discrete codes as in the example of Figure 2. For comparison, we consider two other alternatives. As a baseline approach, we use the Mean Readout utilizing the GNN embeddings of the trained Forge model. The Mean Readout method generates the MIP embedding of an instance by averaging all node features from the GNN hidden layer representation from Figure 1(B). This baseline also behaves as an ablation for Forge without vector quantization. Given the weakness of GNN embeddings in capturing global structure (Feng et al., 2025), we expect this method to perform poorly. Alternatively, starting from the 10-dimensional static node features of the bipartite graph (Fig. 1(A)), we perform two-hop label propagation (Zhu & Ghahramani, 2002) and average the resulting node vectors. Note that global, static instance descriptors (e.g., (Kadioglu et al., 2010)) cannot be used since within D-MIPLIB, every instance in a category shares identical coarse statistics (number of variables, constraints, etc.), rendering them non-informative.
@@
Clustering Visualization: Figure 3 visualizes these embeddings vectors, projected into two dimensions using PaCMAP (Wang et al., 2021). Each dot represents an instances colored by its problem type. As expected, the Mean Readout method looses the global structure leading to an arbitrary clustering. The Label Propagation performs much better capturing some structure, while some problem domains and sizes are mixed in the clustering. Forge embeddings performs the best where each problem type, such as Set Cover, Combinatorial Auction etc., are separated cleanly. Moreover, within each problem domain, the difference in problem difficulty is also observed from the easy to medium and hard categories. Recall again that our Forge model is not trained on these instances. We are able to train on MIPLIB instances and capture the structure of unseen D-MIPLIB instances.
@@
Clustering Accuracy: For quantitative evaluation, we run k-means clustering with 21 clusters, expecting one cluster for each problem type. We then calculate the normalized mutual information score (NMI) between then ground truth problem type assignment vs. k-means clusters generated by each method. The NMI scores are averaged over 10 runs of k-means clustering (the higher, the better). If the predicted clusters are identical to the original problem types, then NMI score would be 1.0, and 0.0 otherwise.
@@
As qualitatively observed from the 2D visualization, the Mean Readout performs poorly with an NMI score of 0.087. This is only slightly better than a uniform random cluster prediction with $1/21=0.047$ NMI score. This is likely due to over-smoothing of averaging dense GNN embedding vectors across a large number of constraints and variables. In contrast, the Label Propagation performs significantly better with an NMI score of 0.790. This method operates directly on the sparse input features of the bipartite graph, thereby avoiding the over-smoothing problem encountered in deep GNN layers. The best performance is achieved by Forge with an NMI score of 0.843. Interestingly, the mean readout embeddings, which perform very poorly, are generated from the same trained GNN within Forge. By utilizing the distribution of the discrete codes of constraints and variables, Forge circumvents the over-smoothing issue and captures the global structure of unseen MIP instances.
@@ 5 Experiments
With this motivation, we study two different downstream tasks: predicting the integrality gap of a given MIP instance, as also used in (Li et al., 2025), and predicting variables for warm-starts, as also used in (Han et al., 2023), in Sections § 5.2 and § 5.3 respectively. Both tasks serve as primal heuristics to enhance MIP solving by obtaining better solutions faster. The integrality gap is used to generate a pseudo-cut added to the original problem formulation to tighten its bound. The warm-start variables are used to provide hints to the solver to start the solving process from a promising initial assignment.
The critical aspect of these sets of experiments is that we use the same pre-trained Forge model to obtain general-purpose MIP embeddings that are then further fine-tuned on a small number of labeled data to learn prediction heads for completely different tasks. This is shown in Figure 4(A). An analogy from NLP would be to generate pre-trained word embeddings using a foundational model, and then to fine-tune prediction heads for entity extraction in a specific domain, e.g., finance, using a small set of labeled data. Our goal is to replicate the success of foundational models in NLP and CV in the context CO.
@@ 5.2 Integrality Gap Prediction & Pseudo-Cut Generation
Integrality Gap Prediction: As covered in the Background Section (§ 2), the integrality gap measures the ratio between the optimal solution of the LP relaxation and the optimal solution of the original integer program. Intuitively, this gap quantifies the quality of the approximation offered by the LP relaxation. A smaller gap means the LP relaxation is a good approximation and is close to value of the best integral solution. Here, we are interested in predicting the integrality gap of a given MIP instance, without having access to or solving for its best integral solution value, as also studied in (Li et al., 2025).
@@
Training Instances: We do not expect this task to generalize between problem classes and/or sets of varying complexities of a given problem class. For instance, there is no reason for an LP gap of 70% to be the same between easy vs. hard Set Cover instances. Figure 4(B) shows the distribution of integrality gap across different problem types and difficulties. As we can see, the integrality gap can be anywhere from 5% to 95% with wide distributions in between. As such, there is no magic constant that one could use at all times heuristically, which makes integrality gap prediction a deliberate learning task. For training, we consider Combinatorial Auction (very-easy, easy, medium), Set Cover (easy, medium, hard), and Generalized Independent Set (easy, medium, hard) with 50 instances for each. In total, we obtain a total of 450 training instances. Note that this is a considerably smaller training set than initially used for training Forge embeddings.
Label Collection: To generate training labels, each training instance is solved using Gurobi with a time limit of 120 seconds. Notice that, when creating this supervised dataset, we do not require solving the instance to optimality which can be computationally challenging. The numeric label is defined as the ratio (or gap) between the integer program solution at the time-out value and the LP relaxation solution. As mentioned, overestimating the integrality gap (or underestimating for maximization problems) can lead to suboptimal solutions, as the solver may terminate prematurely. In contrast, underestimating the gap is generally acceptable, as it still facilitates faster solve times without compromising solution quality. To account for this asymmetry, we adopt a conservative labeling strategy by setting a 120-second timeout when collecting labels. This often results in an underestimation of the true integrality gap, particularly for more challenging instances. By doing so, we reduce the risk of the model overestimating the gap during inference, thereby improving the reliability of the predicted cuts.
Supervised Fine-Tuning: Given the labeled data, to predict the integrality gap, a dense prediction head is added to pre-trained Forge that takes as input the codewords assigned to each node and outputs a real number (Figure 4(A)). The mean readout across all nodes is used as the predicted integrality gap. As a regression task, this is trained with the mean absolute error loss in an end-to-end manner.

### v2

@@ 1 Introduction
With this vision in mind, we propose Forge: Foundational Optimization Representations from Graph Embeddings. Forge is a foundational model to generate useful MIP embeddings through a pre-training framework that learns structural representations at the instance level from a broad distribution of MIP instances without requiring access to their solutions.
@@
  • For the unsupervised setting, we show that our instance level MIP embeddings clusters previously unseen instances across multiple problem types with high accuracy.
  • For the supervised setting, we show that pre-trained Forge embeddings can be fine-tuned for various downstream tasks using a small number of labeled data. We evaluate Forge on two radically different tasks to showcase its versatility: predicting variables for warm-starts and predicting integrality gap for cut-generation. Notably, a single pre-trained Forge model is fine-tuned and applied across diverse problem types including Combinatorial Auctions (CA), Set Covering (SC), Generalized Independent Set (GIS), and Minimum Vertex Cover (MVC).
@@
The remainder of the paper is structured as follows: we start with a brief background of CO in Section §2 followed by the Forge methodology in Section §3. Sections §4 and §5 show experiments to demonstrate the efficacy of Forge in unsupervised and supervised settings respectively.
@@ 3 Forge: Unsupervised Approach to Learn MIP Embeddings
MIP instances are typically represented as a bipartite graph between variables and constraints augmented with node features (Gasse et al., 2019; Ferber et al., 2022; Yau et al., 2024; Chen et al., 2023). This is then followed by training a GNN in a supervised fashion for a specific downstream task on a certain problem class. For example, in Han et al. (2023), the GNN is used for predicting variables for warm-starts trained on Set Cover (SC) and Independent Set (IS) problems. Similarly, in Li et al. (2025), a GNN is used for predicting the integrality gap. Numerous variants follow this template (see §6). Notice that, all of these methods require supervision and do not yield a general-purpose MIP embeddings at the instance level. Taking this a step further; our goal is to learn the structure of MIP instances in an unsupervised manner.
@@
A) MIP-to-BP: Given a MIP instance, we start with its bipartite (BP) representation and node features. Each node in this bipartite graph represents a constraint or a variable, with edges indicating which variables are part of which constraints. Each node is also associated with node features and each edge is weighted by the coefficient of the variable in the constraint. Node features are typically extracted from the internal branch-and-bound search tree when solving an instance. For example, there are 18 input node features used in Gasse et al. (2019). We do not attempt to solve or depend on the solution of the instance. Instead, Forge only uses the basic properties of the input instance. For each constraint node, we introduce 4 features composed of its sense (i.e., >>, << or ==) and the RHS value. For each variable node, we introduce 6 features composed of its type (integer, binary, continuous), upper/lower bound, and the coefficient in the objective function. In total, for each node, we obtain a vector of size 10, padded with zeros accordingly based on node type (Figure 1(A)).
@@
C) Vector Quantized Codebook: To preserve global structure, we introduce a vector quantized codebook with $k$ discrete codes. These codes act as a “vocabulary”, akin to language models, across MIP instances of various domains and difficulties, thereby preserving global structure. The design follows the approaches developed in computer vision (Van Den Oord et al., 2017; Yu et al., 2022; Lee et al., 2022) and the structure-aware graph tokenizer extension in Yang et al. (2024) (Figure 1(C)).
D) GNN-to-CW: GNN embeddings are passed into a vector quantizer which consists of a codebook with $k$ codes. The codebook maps each variable and constraint node to a discrete code. Each code is then mapped into a $d$ dimensional codeword (CW), yielding constraint and variable CW representations. These representations are the same dimension as the hidden GNN layers (Figure 1(D)).
@@
F) Loss Function: The loss function minimizes the edge reconstruction loss, the node feature reconstruction loss and losses related to the vector quantization. Concretely, the loss function is:
@@
Intuitively, the codebook loss in Eq. 3 can be interpreted as k-means clustering, where the codewords $cw_{i}$ (akin to cluster centroids) are moved closer to the node embeddings $h_{i}$ and the node embeddings $h_{i}$ are fixed in place due to the stop-gradient operator. Conversely, the commitment loss in Eq. 4 fixes the codewords $cw_{i}$ using the stop-gradient operator, and instead, moves the embeddings $h_{i}$ towards the codewords. The hyperparameter $\alpha$ weighs the importance of the commitment loss.
@@
Local representations, where each node in the bipartite graph is assigned a discrete code. Each code is further mapped to a codeword which becomes the variable and/or constraint embedding.
Global representations, where we leverage the distribution of codes for the given MIP instance. As shown in Figure 2, each MIP instance is represented with an embedding of size $vocabulary=\lvert codebook\rvert$ where each value denotes the frequency of codes present in the instance. Figure 2 shows a codebook of size 5 (for brevity) and the resulting MIP embedding $\vec{emb}=[3,2,3,2,0]$.
@@ 4 Initial Analysis of Forge Embeddings
Figure 3: 2-D PaCMAP visualization of D-MIPLIB instances using MIP embeddings from (a) the mean readout of node embeddings from the GNN in Forge, (b) the two-hop label propagation of input node features, and (c) the distribution of discrete codes produced by Forge. Normalized Mutual Information (NMI), averaged over 10 runs of $k$-means clustering, is reported for each method.
@@
Training: We use 600 instances from MIPLIB, sorted by size to ensure the resulting bipartite graphs fit on GPU memory. We further generate two more instances from each MIPLIB instance by randomly dropping 5% and 10% of constraints. Note that dropping constraints does not impact feasibility of the instance. In total, we obtain 1,800 MIP instances to train Forge in an unsupervised fashion as described earlier (§3). We use two GraphSage layers with 1024 dimensions for the hidden GNN layers ($d$) and a codebookcodebook with 5000 codes (i.e., the size of the vocabulary).
@@
Testing: We evaluate Forge on 1,050 instances from D-MIPLIB organized into 21 problem type-difficulty pairs. These are distributed as Set Cover (easy, medium, hard), Maximum Independent Set (easy, medium), Minimum Vertex Cover (easy, medium, hard), Generalized Independent Set (easy, medium, hard, very-hard, very-hard2, ext-hard), Combinatorial Auction (very-easy, easy, medium, very-hard, very-hard2), Item Placement (very-hard) and Maritime Inventory Routing Problem (medium), thereby covering a broad spectrum of problem types, sizes and difficulty levels.
@@
MIP Embeddings: These 1,050 instances are passed through Forge to generate one embedding vector per MIP instance using the distribution of discrete codes as shown in Figure 2. For comparison, we consider two other alternatives. As a baseline approach, we use the Mean Readout utilizing the GNN embeddings of the trained Forge model. The Mean Readout method generates the MIP embedding of an instance by averaging all node features from the GNN hidden layer representation from Figure 1(B). This baseline also behaves as an ablation for Forge without vector quantization. Given the weakness of GNN embeddings in capturing global structure (Feng et al., 2025), we expect this method to perform poorly. Alternatively, starting from the 10-dimensional static node features of the bipartite graph (Fig. 1(A)), we perform two-hop label propagation (Zhu & Ghahramani, 2002) and average the resulting node vectors. Note that global, static instance descriptors (e.g., (Kadioglu et al., 2010)) cannot be used since within D-MIPLIB, every instance in a category shares identical coarse statistics (number of variables, constraints, etc.), rendering them non-informative.
@@
Clustering Visualization: Figure 3 visualizes these embeddings vectors, projected into two dimensions using PaCMAP (Wang et al., 2021). Each dot represents an instance colored by its problem type. As expected, the Mean Readout method looses the global structure leading to an arbitrary clustering. The Label Propagation performs much better capturing some structure, while some problem domains and sizes are mixed in the clustering. Forge embeddings performs the best where each problem type, such as Set Cover, Combinatorial Auction etc., are separated cleanly. Moreover, within each problem domain, the difference in problem difficulty is also observed from the easy to medium and hard categories. Recall again that our Forge model is not trained on these instances. We are able to train on MIPLIB instances and capture the structure of unseen D-MIPLIB instances.
@@
Clustering Accuracy: For quantitative evaluation, we run k-means with 21 clusters, expecting one cluster for each problem type. We calculate the normalized mutual information score (NMI) between then ground truth problem type assignment and the k-means clusters generated by each method. The NMI scores are averaged over 10 runs of k-means clustering. If the predicted clusters are identical to the original problem types, then NMI score would be 1.0, and 0.0 otherwise.
@@
As qualitatively observed from the 2D visualization, the Mean Readout performs poorly with an NMI score of 0.087. This is only slightly better than random cluster predictions with an NMI score of $1/21=0.047$. This is likely due to over-smoothing due to averaging dense GNN embedding vectors across a large number of constraints and variables. In contrast, the Label Propagation performs better with an NMI score of 0.790. This method operates directly on the sparse input features of the bipartite graph, thereby avoiding the over-smoothing problem encountered in deep GNN layers. The best performance is achieved by Forge with an NMI score of 0.843. Interestingly, the mean readout embeddings, which perform very poorly, are generated from the same trained GNN within Forge. By utilizing the distribution of the discrete codes of constraints and variables, Forge circumvents the over-smoothing issue and captures the global structure of unseen MIP instances.
@@ 5 Experiments
We therefore study two different downstream tasks: predicting the integrality gap of a given MIP instance, as also used in (Li et al., 2025), and predicting variables for warm-starts, as also used in (Han et al., 2023), in Sections § 5.2 and § 5.3 respectively. Both tasks serve as primal heuristics to enhance MIP solving by obtaining better solutions faster. The integrality gap is used to generate a pseudo-cut added to the original problem formulation to tighten its bound. The warm-start variables are used to provide hints to the solver to start the solving process from a promising initial assignment.
A critical aspect of these experiments is that we use the same pre-trained Forge model to obtain general-purpose MIP embeddings that are then fine-tuned on a small number of labeled data to learn prediction heads for completely different tasks (as shown in Figure 4(A)). An analogy from NLP would be to generate pre-trained word embeddings using a foundational model, and then fine-tune prediction heads for entity extraction in a specific domain, e.g., finance, using a small set of labeled data. Our goal is to replicate the success of foundational models in NLP and CV in the context CO.
@@ 5.2 Integrality Gap Prediction & Pseudo-Cut Generation
Integrality Gap Prediction: As covered earlier (§ 2), the integrality gap measures the ratio between the optimal solution of the LP relaxation and the optimal solution of the original integer program. Intuitively, this gap quantifies the quality of the approximation offered by the LP relaxation. A smaller gap means the LP relaxation is a good approximation and is close to the value of the best integral solution. Here, we are interested in predicting the integrality gap of a given MIP instance, without solving for its optimal integral solution value, as also studied in (Li et al., 2025).
@@
Training Instances: We do not expect this task to generalize between problem classes and/or sets of varying complexities of a given problem class. For instance, there is no reason for an LP gap of 70% to be the same between easy vs. hard SC instances. Figure 4(B) shows the distribution of integrality gap across different problem types and difficulties. Note that the integrality gap can be anywhere from 5% to 95% with wide distributions in between. As such, there is no magic constant that one could use at all times heuristically, which makes integrality gap prediction a deliberate learning task. For training, we consider CA (very-easy, easy, medium), SC (easy, medium, hard), and GIS (easy, medium, hard) with 50 instances for each. In total, we obtain a total of 450 training instances. Note that this is a considerably smaller training set than initially used for training Forge embeddings.
Label Collection: To generate training labels, each training instance is solved using Gurobi with a time limit of 120 seconds. Note that this does not require solving the instance to optimality which can be computationally challenging. The numeric label is defined as the ratio (or gap) between the integer program solution at the time-out value and the LP relaxation solution. As mentioned, overestimating the integrality gap (or underestimating for maximization problems) can lead to suboptimal solutions, as the solver may terminate prematurely. In contrast, underestimating the gap is generally acceptable, as it still facilitates faster solve times without compromising solution quality. This asymmetry is accounted for by adopting a conservative labeling strategy by setting a 120-second timeout when collecting labels. This often results in an underestimation of the true integrality gap, particularly for more challenging instances. By doing so, we reduce the risk of the model overestimating the gap, thereby improving the reliability of the predicted cuts.
Supervised Fine-Tuning: Given the labeled data, a dense prediction head is added to the pre-trained Forge model that takes as input the codewords assigned to each node and outputs a real number (Figure 4(A)). The mean readout across all nodes is used as the predicted integrality gap. As a regression task, this is trained with the mean absolute error loss in an end-to-end manner.